OpenStudy (anonymous):

How many numbers are there b/w 100 and 1000 such that at least one of their Digit is 7?

OpenStudy (anonymous):

@hartnn @eliassaab @Zarkon

OpenStudy (anonymous):

there are 900 three digit numbers b/w 100 and 1000

hartnn (hartnn):

numbers between 100 and 1000 = 900 numbers between 100 and 1000 which do not have digit 7 in any place = 8 x 9 x 9 = 648 Unit digit could take any value of the 9 values (0 to 9, except 7) Tens Digit could take any value of the 9 values (0 to 9, except 7) Hundreds digit could take any value of the 8 values (1 to 9, except 7) therefore numbers between 100 and 1000 which have atleast one digit as 7 = 900 - 648 = 252

OpenStudy (anonymous):

999 = 100 + n -1 n = 900

OpenStudy (anonymous):

What about this 035 @hartnn u counted this also

OpenStudy (anonymous):

thxxx...

hartnn (hartnn):

nopes,i didn't,thats why there are 8 instead of 9 values at hundred's place.

OpenStudy (anonymous):

oh. got it. @hartnn is correct

OpenStudy (anonymous):

There are 252 here they are: {107, 117, 127, 137, 147, 157, 167, 170, 171, 172, 173, 174, 175, \ 176, 177, 178, 179, 187, 197, 207, 217, 227, 237, 247, 257, 267, 270, \ 271, 272, 273, 274, 275, 276, 277, 278, 279, 287, 297, 307, 317, 327, \ 337, 347, 357, 367, 370, 371, 372, 373, 374, 375, 376, 377, 378, 379, \ 387, 397, 407, 417, 427, 437, 447, 457, 467, 470, 471, 472, 473, 474, \ 475, 476, 477, 478, 479, 487, 497, 507, 517, 527, 537, 547, 557, 567, \ 570, 571, 572, 573, 574, 575, 576, 577, 578, 579, 587, 597, 607, 617, \ 627, 637, 647, 657, 667, 670, 671, 672, 673, 674, 675, 676, 677, 678, \ 679, 687, 697, 700, 701, 702, 703, 704, 705, 706, 707, 708, 709, 710, \ 711, 712, 713, 714, 715, 716, 717, 718, 719, 720, 721, 722, 723, 724, \ 725, 726, 727, 728, 729, 730, 731, 732, 733, 734, 735, 736, 737, 738, \ 739, 740, 741, 742, 743, 744, 745, 746, 747, 748, 749, 750, 751, 752, \ 753, 754, 755, 756, 757, 758, 759, 760, 761, 762, 763, 764, 765, 766, \ 767, 768, 769, 770, 771, 772, 773, 774, 775, 776, 777, 778, 779, 780, \ 781, 782, 783, 784, 785, 786, 787, 788, 789, 790, 791, 792, 793, 794, \ 795, 796, 797, 798, 799, 807, 817, 827, 837, 847, 857, 867, 870, 871, \ 872, 873, 874, 875, 876, 877, 878, 879, 887, 897, 907, 917, 927, 937, \ 947, 957, 967, 970, 971, 972, 973, 974, 975, 976, 977, 978, 979, 987, \ 997}

OpenStudy (anonymous):

lol......thxxx....@eliassaab plzz luk my new question http://openstudy.com/study?signup#/updates/5044bd51e4b0ff4ca7b7e9d2